(1/4,-1/2)& (3/4,4)
what is the equation of the line?
y=

Found 2 solutions by texttutoring, solver91311:
Answer by texttutoring(324)

Use the slope equation.

m=(y2-y1)/(x2-x1) with x1=1/4, y1=-1/2 and x2=3/4, y2=4

m = (4-(-1/2)) / (3/4 - 1/4)
m = (9/2) / (2/4)
m = (9/2)*(4/2)
m = 36/4
m=9

Now use y=mx+b, with y=4, x=3/4 and m = 9. Solve for b.

y=mx+b
4 = 9(3/4) + b
4 = 27/4 + b
4 - 27/4 = b
16/4 - 27/4 = b
-11/4 = b

The equation of the line is:

y = 9x -11/4

Answer by solver91311(24713)


Use the two-point form of the equation of a line:



where and are the coordinates of the two given points.

Once you substitute the given values and simplify by performing any indicated arithmetic, you will have an equation of the desired line. You cannot find the equation of any line because any given line has an infinite number of equivalent representations. You did not specify what form your answer needs to take, so you need to check with your instructor to determine if the simplfied two-point form is acceptable, or if you need to put your equation into another form, typically slope-intercept () or standard form, (), and then rearrange your equation appropriately.

John